OpenStudy (anonymous):

23 HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! 3 + a + b = a + 3 + b which algebraic property is this

OpenStudy (jiteshmeghwal9):

commutative property :)

OpenStudy (jiteshmeghwal9):

\[\Huge{\color{red}{commutative \space property:}}\]\[\Huge{\color{green}{a+b}}=\color{green}{b+a}\]

OpenStudy (anonymous):

what about Solve: - 1 over 2 x + 3 = −x + 7